If there are 0.0345 mols of HI in 47 mL of solution, what is its concentration? Also, what
is the pH?
1
Expert's answer
2020-05-06T13:59:16-0400

1.C(HI)=n(HI)/V=0.0345/0.047=0.734 mol/L

2.HI<=>H++I-;

pH=-lg[H+]=-lg(0.734)=0.134